Jump to main content.


Drake Chemical

EPA Weekly Situation Reports for July 1998

U. S. ARMY CORPS OF ENGINEERS
DRAKE CHEMICAL SUPERFUND SITE

182 MYRTLE STREET
LOCK HAVEN, PA 17745
PHONE 717-748-4505
FAX 717-748-7549

Wednesday, July 29, 1998

FOR DATES: From 7:00 a.m. Wednesday, July 22, 1998 to 7:00 a.m. Wednesday, July 29, 1998.

The Thermal Destruction Facility (TDF) was on-line for 103 hours and processed 3424.4 tons of contaminated soil during the time frame of this report. The TDF was on-line 61.3 % of the time with an average hourly feed rate of 33.2 tons per hour while on-line. The contractor continues to work 24 hours a day, 7 days a week operating and maintaining the TDF.

The contractor has successfully treated approximately 94,020 tons of contaminated soil through the TDF amounting to 34.1 % of the overall contaminated soil on-site.

The TDF was placed in a cold shutdown mode on Sunday, July 16, in order to perform interlock checks and maintenance to the interior components of the TDF. Feed was stopped at approximately 1836 hours on Sunday evening in preparation for the interlock checks. A planned Thermal Relief Valve (TRV) opening was performed at 2130 hours Sunday in order to test the TDF's interlock shut down system. The TRV was reopened at 0740 hours on Monday morning for an extended maintenance period.

During the cold shut down, the contractor's main focus will be to replace the kiln feed screws, the feed screw cooling jacket, and to perform preventative maintenance on the Induced Draft (ID) fan motor. The contractor's TDF operations crew will also be performing maintenance on several other important components of the TDF. The contractor plans to begin heating up the kiln as early as Saturday, August 1, 1998.

For the period of time prior to the cold shut down, the TDF was on-line 95.3% of the time.

Construction activities during the time frame of this report included wastewater management, excavation of soils within the exclusion zone, feed preparation, debris separation, lime treatment of soils, and rock and concrete debris handling and backfilling. The contractor continued with real time perimeter air monitoring and off-site air monitoring activities.

Related site activities during this period included:

* Weekly progress meeting held on Tuesday, July 28, 1998. * MG Jerry Sinn, and his USACE North Atlantic Division Staff, toured the Drake Chemical Superfund Site with high-ranking representatives from USACE, OHM-IT, and PADEP. The General was able to observe one of the Corps' largest Superfund sites, and witness the partnering between the contractor and government agencies.

Major activities for the next week include: Restarting and operation of the TDF, soil feed preparation, excavation, off-site and perimeter air monitoring activities, WWTP operations, ash handling, ash backfilling, rock & inorganic debris handling, rock cleaning & backfilling, drum handling, and drum bulking.

PREPARED BY: Steve Gibson for David Modricker, USACE


Wednesday, July 22, 1998

FOR DATES: From 7:00 a.m. Wednesday, July 15, 1998 to 7:00 a.m. Wednesday, July 22, 1998.

The Thermal Destruction Facility (TDF) was on-line for 161.64 hours and processed 6127.9 tons of contaminated soil during the time frame of this report. The TDF was on-line 96.2% of the time with an average hourly feed rate of 37.9 tons per hour while on-line. The contractor continues to work 24 hours a day, 7 days a week operating the TDF.

The contractor has successfully treated approximately 87,334.6 tons of contaminated soil through the TDF amounting to 31.6% of the overall contaminated soil on-site.

During the time frame of this report period, feed was stopped several times to remove soil build-ups on the feed screws and to perform daily calibrations on the TDF. Feed was manually stopped several times as a precautionary measure to possible TDF power failure during local lightning storm events. Feed stoppages resulted in only 6.4 hours of down time for the entire week.

OHM and the U.S. Army Corps of Engineers continue to evaluate the modified heat stress monitoring program, which was implemented to more carefully monitor workers during periods of extreme heat. Since the modifications have been made to the heat stress monitoring program, no incidents of heat exhaustion have been reported by site workers.

A team of inspectors, led by OSHA, completed a four day audit of the health and safety program and work practices utilized at the Drake Chemical Superfund Site. The inspection was part of a national program developed by USEPA and OSHA to inspect hazardous waste Superfund sites utilizing incineration. At the conclusion of the inspection on Thursday afternoon, the inspection team held an exit-briefing meeting with all relevant site representatives. The field observations that were noted were corrected immediately. The final audit report was to be forwarded in four to six weeks.

Construction activities during the time frame of this report included wastewater management, the restarting of excavation of soils within the exclusion zone, feed preparation, debris separation, lime treatment of soils, and rock and concrete debris handling, cleaning and backfilling. The contractor continued with real time perimeter air monitoring and off-site air monitoring activities.

Related site activities during this period included: * Weekly progress meeting was held on Tuesday, July 21, 1998. * The Drake Site family picnic was held on Sunday, July 19, 1998. Site tours were given to friends and family of the employees working at Drake. * A field manager partnering meeting between OHM, USACE, and PADEP was held on July 17, 1998 to discuss concerns and topics relevant to the Drake Chemical Superfund Site .

There is a scheduled TDF cold shutdown planned to start Sunday, July 26, 1998 and to last approximately one week. While the unit is shutdown, preventative maintenance is to be performed on the induced draft (ID) fan, the feed screws are to be replaced and general maintenance on the TDF is to be performed. During this same time period soil feed preparation, off-site and perimeter air monitoring activities, WWTP operations, ash handling, ash backfilling, rock & inorganic debris handling, drum handling, and drum bulking are to continue.

PREPARED BY: Steve Gibson for David Modricker, USACE


Wednesday, July 15, 1998

FOR DATES: From 7:00 a.m. Wednesday, July 8, 1998 to 7:00 a.m. Wednesday, July 15, 1998.

The Thermal Destruction Facility (TDF) was on-line for 164.8 hours and processed 6635.4 tons of contaminated soil during the time frame of this report. The TDF was on-line 98.1 % of the time with an average hourly feed rate of 40.3 tons per hour while on-line. The contractor continues to work 24 hours a day, 7 days a week operating the TDF.

The contractor has successfully treated approximately 82,557 tons of contaminated soil through the TDF amounting to 29.9 % of the overall contaminated soil on-site.

During the time frame of this report period, feed was stopped several times to perform routine maintenance and calibrations on the TDF. Feed stoppages resulted in only 3.2 hours of down time for the entire week.

During the past week, two bins of treated soil failed the laydown criteria for backfilling on site and the material was returned to the exclusion zone for additional thermal treatment. All other treated soil during the week passed laydown criteria.

OHM and the U.S. Army Corps of Engineers continue to evaluate the modified heat stress monitoring program, which was implemented to more carefully monitor workers during periods of extreme heat. Since the modifications had been made to the heat stress monitoring program, no incidents of heat exhaustion have been reported by site workers.

Construction activities during the time frame of this report included wastewater management, excavation of soils within the exclusion zone, feed preparation, debris separation, lime treatment of soils, and rock and concrete debris handling and backfilling. The contractor continued with real time perimeter air monitoring and off-site air monitoring activites.

New excavation of contaminated soil from the south side of the stockpile was not accomplished during the time frame of this report.

Related site activities during this period included:

* Weekly progress meeting held on Tuesday, July 14, 1998. * The third monthly stack testing and sampling event for dioxins and furans was completed on July 14, 1998. * A team of inspectors, led by OSHA, is on site performing an audit of the health and safety program and work practices. The inspection is part of a national program developed by USEPA and OSHA to inspect hazardous waste Superfund sites utilizing incineration.

Major activities for the next week include: Operating the TDF, soil feed preparation, off-site and perimeter air monitoring activities, WWTP operations, ash handling, ash backfilling, rock & inorganic debris handling, drum handling, and drum bulking.

PREPARED BY: Danny Zion for David Modricker, USACE


Wednesday, July 8, 1998

FOR DATES: From 7:00 a.m. Wednesday, July 1, 1998 to 7:00 a.m. Wednesday, July 8, 1998.

The Thermal Destruction Facility (TDF) was on-line for 165.7 hours and processed 7013 tons of contaminated soil during the time frame of this report. The TDF was on-line 98.6 % of the time with an average hourly feed rate of 42.3 tons per hour while on-line. The contractor continues to work 24 hours a day, 7 days a week operating the TDF.

The contractor has successfully treated approximately 76,672.9 tons of contaminated soil through the TDF amounting to 27.8 % of the overall contaminated soil on-site.

During the time frame of this report period, feed was stopped several times to perform routine maintenance and calibrations on the TDF. Feed stoppages lasted only a few minutes at each event and resulted in only 2.3 hours of down time for the entire week.

OHM and the U.S. Army Corps of Engineers continue to evaluate the modified heat stress monitoring program, which was implemented to more carefully monitor workers during periods of extreme heat. Since the modifications had been made to the heat stress monitoring program, no incidents of heat exhaustion have been reported by site workers.

Construction activities during the time frame of this report included wastewater management, excavation of soils within the exclusion zone, feed preparation, debris separation, lime treatment of soils, and rock and concrete debris backfilling on the bottom 2 feet of the excavation. The contractor continued the placement of Concover on the stockpile and placement of odor suppressant foam on the excavation face within the exclusion zone. The contractor continued the real time perimeter air monitoring and the off-site air monitoring activites.

Excavation within the exclusion zone uncovered a large amount of contaminated demolition debris and drum carcasses. The demolition debris will be separated and the drum carcasses will be shipped to a hazardous waste landfill. No intact drums were reported to be uncovered during site excavation operations this week.

Related site activities during this period included:

* Weekly progress meeting held on Tuesday, July 7, 1998. * The third monthly stack testing and sampling event for dioxins and furans is scheduled for July 14 through July 15, 1998. * The City of Lock Haven Environmental Advisory Committee held their monthly meeting on July 7, 1998 at the Drake Chemical Superfund Site PR Trailer. USACE and USEPA personnel later gave the committee tour of the site.

Major activities for the next week include: Operating the TDF, soil feed preparation, off-site and perimeter air monitoring activities, WWTP operations, ash handling, ash backfilling, rock & inorganic debris handling, drum handling, and drum bulking.

PREPARED BY: Steve Gibson & David Modricker, USACE


Wednesday, July 1, 1998

FOR DATES: From 7:00 a.m. Wednesday, June 24, 1998 to 7:00 a.m. Wednesday, July 1, 1998.

The Thermal Destruction Facility (TDF) was on-line for 103.8 hours and processed 4344.8 tons of contaminated soil during the time frame of this report. The TDF was on-line 61.8 % of the time with an average hourly feed rate of 41.9 tons per hour while on-line. The contractor continues to work 24 hours a day, 7 days a week operating the TDF.

To date, the contractor has successfully treated approximately 70,159 tons of contaminated soil through the TDF amounting to 25 % of the overall contaminated soil on-site.

At the beginning of this reporting period, the TDF remained in hot standby while the soil feed screws and the feed screw cooling jacket were replaced. Hazardous waste feed resumed at approximately 1015 hours on Friday, June 26, 1998.

During the remainder of this report period, feed was stopped several times to perform routine maintenance on the newly installed feed screws. Feed was also manually stopped several times as a precaution during local lightning storm activity. The contractor suspended feed to the incinerator on Monday morning, June 29, to evaluate their heat stress monitoring program. Hazardous waste feed to the incinerator resumed Monday evening. OHM implemented additional measures to monitor workers for signs of heat stress. OHM and the U.S. Army Corps of Engineers will continue to evaluate the new heat stress monitoring program as work progresses during the summer months.

During the hot standby period, an unscheduled Thermal Relief Valve (TRV) opening occurred as a result of a tripped circuit breaker in the control room during TDF maintenance. Feed to the incinerator had been shut down for over 40 hours prior to the opening. The contractor made the proper notifications and the TRV was closed approximately 5 minutes later.

Construction activities during this period included wastewater management, excavation of soils within the exclusion zone, rock washing and decontamination of inorganic debris for backfilling, bulking of compatible drums, and overpacking of previously excavated drums. The contractor continued the placement of Concover on the stockpile and placement of odor suppressant foam on the excavation face within the exclusion zone, and continued off-site and perimeter air monitoring activities.

Excavation within the exclusion zone uncovered a large amount of contaminated demolition debris and drum carcasses. The demolition debris will be separated and the drum carcasses will be shipped off-site for disposal. Drum bulking activities began during the time frame of this report. The contents of ten overpacked drums were combined in accordance with chemical compatibility in preparation for shipment off-site for disposal. Drum bulking activities were later postponed due to high temperatures during the day. Bulking activities may be continued during the cooler hours at night.

Related site activities during this period included:

* Weekly progress meeting held on Tuesday, June 30, 1998. * The USACE, USEPA, PADEP, OHM, and the City of Lock Haven met to discuss the final grading plan for the site.

Major activities for the next week include: Operating the TDF, soil feed preparation, off-site and perimeter air monitoring, WWTP operations, ash handling, ash backfilling, rock & inorganic debris handling, drum handling, and drum bulking.

PREPARED BY: Steve Gibson & David Modricker, USACE

Region 3 | Mid-Atlantic Cleanup | Mid-Atlantic Superfund |EPA Home | EPA Superfund Homepage


Local Navigation


Jump to main content.